OpenStudy (anonymous):

log4(16x^2y^5/z^4t^8) rewrite in terms of multiple logarithms of single variables

OpenStudy (anonymous):

\[\log_{4}16+\log_{4}x^2+\log_{4}y^5-\log_4{z^4}-\log_{4}t^8\] \[=2+2\log_{4}x+5\log_{4}y-4\log_{4}z-8\log_{4}t\]

OpenStudy (anonymous):

\[\log_aMN=\log_aM+\log_aN\] \[\log_a\frac{M}{N}=\log_aM-\log_aN\]

OpenStudy (anonymous):

thanks dock can you rewrite this as a single logarithm 2ln(4)-4ln(x)+1/2ln(y)-4ln(z)+6ln(w)+3ln(t)

OpenStudy (anonymous):

use the rules above

OpenStudy (anonymous):

\[\ln(x)=\log_e(x)\]

OpenStudy (anonymous):

is it ln(16 * sqrty*x^4/z^4*w^6*t^3)

OpenStudy (anonymous):

x^4 should be in the denominator

OpenStudy (anonymous):

w^6 in the numerator

OpenStudy (anonymous):

t^3 in the numerator

OpenStudy (anonymous):

it would be easier if you arranged it so all of the addition is on the left, all of the subtraction is on the right